Calculate harmonic mean of the variable x from the following data.

U = X-3.5/0.5 -3 -2 -1 0 1 2 3


1
Expert's answer
2021-08-12T12:54:17-0400

Solution:

Given data: -3.5/0.5, -3, -2, -1, 0, 1, 2, 3

Or -7, -3, -2, -1, 0, 1, 2, 3

Here, n = 8

Now, harmonic mean"=\dfrac{n}{\dfrac{1}{x_1}+\dfrac{1}{x_2}+...+\dfrac{1}{x_n}}"

"=\dfrac{8}{\dfrac{1}{-7}+\dfrac{1}{-3}+\dfrac{1}{-2}+\dfrac{1}{-1}+\dfrac{1}{0}+\dfrac{1}{1}+\dfrac{1}{2}+\dfrac{1}{3}}
\\=\dfrac{8}{\infty} \ [\because \frac10=\infty]
\\=0"